If you pay $55 for a share of common stock that has a constant growth rate of 6% and it is expected to pay a dividend of $1.25 what would be your return?
Answer #1

Given,

The price of the stock, P0 = $55.

Growth rate, g = 6% or 0.06.

Dividend expected, D1 = $1.25

Let, expected rate of return, r*

We know that, P0 = D1 / ( r* - g )

=> ( r* - g ) = D1 / P0 = 1.25/55 = 0.0227

=> r* = 0.0227 + g = 0.0227 + 0 06 = 0.0827

Therefore, r* = 0.0827 or 8.27%